OpenStudy (andijo76):

This is a long shot but I need help if anyone here can help budget(x) 63,87,49,31,193,96,91 gross(y) 61,68,45,62,536,142,48 budget (x) =63,87,49,31,193,96,91=610 gross(y)= 61,68,45,62,536,142,48=962 x*y= 3843,5916,2205,1922,103448,13632,4368=130966 x^2= 3962,7569,2401,961,37249,9216,8281=70608 y^2= 3721,4624,2025,3844,287296,20164,2304=323978 i will show what i have so far

OpenStudy (andijo76):

numerator of r=(130966)-(610)(962)/7= 47134.571428 denominator of r= SQRT [7068-(610)^2/7]*SQRT[323978(962)^2/7] =Sqrt[17450.857142] =Sqrt[191771.714285] r=17450.857142 now I am stuck

OpenStudy (andijo76):

there is a question i will post

OpenStudy (andijo76):

brb

OpenStudy (andijo76):

Listed below are the budgets (in millions of dollars and the gross receipts (in millions of dollars) for randomly selected movies construct a scatterplot find the value of the linear correlation coefficient r and find the P value using a=0.05 is there sufficient evidence to conclude that there is a linear correlation between budgets and gross receipts do the results change if the actual budgets listed are $63,000.000, $87,000.000, 49,000.000 and so on

OpenStudy (andijo76):

r is wrong

OpenStudy (the_fizicx99):

Ugh I hated statistics a few assignments back >_< let me try something, this is the formula I had to use to find R (Correlation Coefficient)

OpenStudy (the_fizicx99):

I plugged in your x and y values and got this

OpenStudy (the_fizicx99):

And this is what your residual plot should look like if you're going to use one,

OpenStudy (andijo76):

r = ? / [ ? ] * [ ? ] this was the next step what do i put here i think the second number is [132.10169 and the second number is [437.91747]

OpenStudy (the_fizicx99):

I used a program that calculates it for you. The formula itself is quite cumbersome if you mess up on just one number -- happened to me. You can try to do it by hand: n = Number of values or elements x = First variable y = Second variable Σxy = Sum of the product of first and second variables Σx = Sum of first variable Σy = Sum of second variable Σx^2 = Sum of the square of the first variable Σy^2 = Sum of the square of the second variable So simply first find the values and plug in into the formula. Simplifying isn't that hard part, though finding them is somewhat difficult >_<

OpenStudy (andijo76):

what is the test statistic t

OpenStudy (the_fizicx99):

what?

OpenStudy (andijo76):

that is what is i asking for the test statistic t round to three decimal places as needed this is the first i have seen this

OpenStudy (the_fizicx99):

So have I, yeah I've never been taught that lol, I don't want to misguide you o.o

OpenStudy (andijo76):

i found a website i typed it in and it gave it to me

OpenStudy (the_fizicx99):

ah ok, nice :p

OpenStudy (andijo76):

where did you get the program you used

OpenStudy (the_fizicx99):

My school advised to use it, its called geogebra. http://www.geogebra.org/cms/en/ It takes a while to get used to

OpenStudy (andijo76):

thanks i can use all the help i can get and thanks for your help

OpenStudy (the_fizicx99):

Yw :P